### 28 Feb 2026 v1

# Theory of Code Space: Do Code Agents Understand Software Architecture?

Abstract
AI code agents excel at isolated tasks—writing functions, fixing localized bugs—yet struggle with complex, multi-file software engineering that requires understanding how dozens of modules relate to each other. We hypothesize that these failures stem from an inability to construct, maintain, and update coherent architectural beliefs during codebase exploration. We introduce Theory of Code Space (ToCS), a benchmark that evaluates this capability by placing agents in procedurally generated codebases under partial observability, requiring them to build structured belief states over module dependencies, cross-cutting invariants, and design intent. Our framework makes three key design choices: (1) a procedural codebase generator producing medium-complexity Python projects with four typed edge categories reflecting different discovery methods—from syntactic imports to config-driven dynamic wiring—with planted architectural constraints and verified ground truth; (2) a partial observability harness where agents explore under a budget, ensuring that exploration strategy matters; and (3) periodic belief probing via structured JSON, producing a time-series of architectural understanding rather than a single final snapshot. We decompose the Active-Passive Gap from spatial reasoning benchmarks into selection and decision components, and introduce Architectural Constraint Discovery as a code-specific evaluation dimension absent from spatial domains. Preliminary experiments with four rule-based baselines and five frontier LLM agents from three providers validate the benchmark’s discriminative power: methods span a wide performance range (F1 from 0.129 to 0.646), LLM agents discover semantic edge types invisible to all baselines, yet weaker models score below simple heuristics—revealing that belief externalization, faithfully serializing internal understanding into structured JSON, is itself a non-trivial capability and a first-order confounder in any belief-probing benchmark. We release the complete toolkit as open-source software and invite the community to evaluate additional models and contribute codebase patterns.111Code: https://github.com/che-shr-cat/tocs
1 Introduction
Large language models achieve remarkable scores on code generation benchmarks (Chen et al., 2021; Austin et al., 2021), creating an expectation that they possess deep understanding of software architecture. Yet practitioners report a persistent gap: models that solve HumanEval problems with ease produce incoherent results when modifying real codebases with dozens of interdependent modules (Jimenez et al., 2024). This gap has resisted satisfying explanation.
Recent work on spatial reasoning offers a compelling diagnostic framework. Zhang et al. (2026) introduced the Theory of Space (ToS) benchmark, demonstrating that multimodal models fail to maintain coherent internal “cognitive maps” when actively exploring partially observable environments. They identified two core phenomena: an Active-Passive Gap (models degrade when they must gather information themselves rather than receiving it upfront) and Belief Inertia (models cannot update their spatial beliefs when the environment changes). We hypothesize that the same latent state maintenance failures explain the struggles of code agents. The term “cognitive map” traces to Tolman (1948); in program comprehension, Pennington (1987) and von Mayrhauser and Vans (1995) showed that developers build layered mental models of code—our benchmark operationalizes and measures this process for AI agents.
We propose Theory of Code Space (ToCS), a benchmark that transplants this diagnostic framework to software engineering. Rather than grid worlds, our “environment” consists of procedurally generated codebases with controlled architectural structure. Rather than allocentric spatial maps, our “cognitive maps” are structured representations of module dependencies, typed edges, cross-cutting invariants, and design intent. The agent explores under partial observability—opening files one at a time under a budget—and must externalize its architectural belief as structured JSON at regular intervals.
Beyond the spatial analogy, we introduce Architectural Constraint Discovery—a dimension absent from spatial benchmarks. In code, architectural decisions encode checkable constraints: a forbidden dependency enforces a service boundary; a validation chain ensures data integrity. We plant explicit, verifiable constraints in our generated codebases and measure whether agents can discover them.
Our contributions are:
1.
The ToCS benchmark framework: the first evaluation for active architectural belief construction, revision, and exploitation in code (§3).
2.
A procedural codebase generator producing medium-complexity projects with four typed edge categories, planted invariants, and verified ground truth (§4).
3.
Pilot experiments with four baselines and five frontier LLM agents from three providers, showing that the strongest agent surpasses all baselines (F1 = 0.646 vs. 0.577) while revealing a wide capability spread: LLM agents discover all four edge types but vary 6 in belief externalization fidelity (§6).
4.
An open benchmark with evaluation harness, scoring pipeline, and community call for expanded model evaluation (§7).
2 Related Work
Code Generation Benchmarks.
SWE-bench (Jimenez et al., 2024) evaluates bug-fixing in real GitHub repositories but measures patch correctness, not evolving architectural understanding. ContextBench (Li et al., 2026) advances this by logging agent trajectories and scoring context retrieval precision/recall against human-verified gold contexts—measuring what the agent looked at but not what the agent believes about architecture. SWE-ContextBench (Zhu et al., 2026) tests cross-task experience reuse. RepoBench (Liu et al., 2024) and LoCoBench-Agent (Qiu et al., 2025) evaluate repo-level completion and long-context interaction respectively, with LoCoBench observing a “comprehension-efficiency trade-off” that we formalize as the Active-Passive Gap. RefactorBench (Masai et al., 2025) targets multi-file refactoring but evaluates output correctness without probing belief state. None of these benchmarks require agents to externalize a revisable architectural belief state, impose exploration budgets, or score against typed dependency graphs with planted constraints. We note that our own belief externalization finding (§6) reveals a complementary limitation: even when agents do externalize beliefs, what they report may not reflect what they internally represent.
Code Understanding Tools.
CodePlan (Bairi et al., 2024) augments LLMs with static-analysis dependency graphs for change propagation—an engineering solution that validates our premise: agents need architectural maps but currently cannot build them alone. Aider’s RepoMap (Gauthier, 2024) uses tree-sitter parsing and PageRank to construct repository context. Code World Models (FAIR CodeGen team et al., 2025) train LLMs on execution traces to improve code reasoning. ToCS provides the diagnostic benchmark to test whether such approaches actually improve architectural belief quality.
Spatial Reasoning.
Theory of Space (Zhang et al., 2026) demonstrated Active-Passive Gap and Belief Inertia in grid-world environments with multimodal models. Related spatial benchmarks include SpatialVLM (Chen et al., 2024), Habitat (Savva et al., 2019), and OpenEQA (Majumdar et al., 2024), which evaluate spatial reasoning and embodied question answering but do not require constructing revisable belief states. We transplant this framework to code, adding architectural constraint discovery (absent from spatial domains where object placement “why” is rarely meaningful). TOM-SWE (Zhou et al., 2025) models the user’s mental state in SWE agents—complementary to our focus on the agent’s belief about the codebase.
3 Theory of Code Space
3.1 Problem Formulation
We formalize codebase understanding as belief construction over a latent architectural state. Let denote the ground-truth architecture, comprising a typed dependency graph with edge types , a set of cross-module invariants , and exported contracts . At time , the agent has history and we evaluate three operations on its belief . Since is a latent internal state, we measure it through probing: periodically asking the agent to externalize its belief as structured JSON (§3.3). We evaluate:
- Construct
Build from partial observations gathered through active exploration.
- Revise
Update when the environment mutates () and the agent encounters evidence of the change.
- Exploit
Use to perform a downstream engineering task correctly (v1.0; v0.1 uses counterfactual probes as a proxy).
3.2 Environment and Action Space
The agent interacts with the codebase through five actions with fixed tool semantics:
LIST(): Filenames in directory (no contents, no recursion).
OPEN(): Full contents of file . Costs 1 action.
SEARCH(): Matching filepaths + line numbers. No content snippets. Costs 1 action.
INSPECT(, ): Type signature + docstring of symbol . No body. Costs 1 action.
DONE(): Terminate. Costs 0 actions.
The agent operates under a budget (default ). Critically, SEARCH returns only locations, never content—ensuring that architectural understanding requires deliberate OPEN decisions.
3.3 Cognitive Map Probing
Every actions, the harness interrupts the agent with a structured prompt requesting it to externalize its current architectural belief as JSON. Critically, probing is free: it does not consume an action from the budget , and the agent retains the same remaining budget before and after a probe. This ensures that measurement does not interfere with exploration strategy—agents are evaluated on their understanding, not on a trade-off between exploring and reporting.
The cognitive map contains: (1) component beliefs with status (observed/inferred/unknown), purpose, exported symbols with typed signatures, and typed dependency edges with per-element confidence; (2) invariant beliefs with structured canonical form (type, src, dst, via, pattern) and evidence pointers; (3) uncertainty tracking—explicit list of unexplored regions. The resulting time-series of maps (one every steps) captures how understanding develops, not just the final state.
3.4 Evaluation Modes
ToCS decomposes the Active-Passive Gap through four conditions:
Active: Agent chooses actions under budget .
Passive-Full: Agent receives the entire codebase; probed once.
Passive-Oracle: Agent receives files selected by oracle (maximum ground-truth connectivity); one file per step, probed every .
Passive-Replay: Agent receives the exact observation trace from a prior active run, without making decisions. Each observation counts as one step.
This decomposes: (passive-full active) into (passive-oracle active, the cost of choosing which files) and (passive-replay active, the cost of deciding what to do with observations).
Scope note: v0.1 evaluates the Construct operation in Active mode only. The passive conditions and APG decomposition are implemented in the harness and reserved for future work with frontier model evaluations.
3.5 Architectural Constraint Discovery
Beyond structural mapping, we probe whether agents discover planted architectural constraints:
Forbidden dependency: “Module must not import module directly.”
Interface-only access: “Module must access only through interface .”
Validation chain: “Data must pass through validation before reaching module .”
Each constraint has a discoverability requirement: test evidence, structural patterns, or documentation in the codebase. Scored via counterfactual multiple-choice probes (“Which change would violate an architectural constraint?”).
4 Procedural Codebase Generation
4.1 Architecture Grammar
Our v0.1 generator produces Pipeline architecture codebases with controlled anti-triviality measures. A PipelineTemplate grammar defines:
Domain pools: Three domains (data ETL, log processing, text processing), each with 8 semantically coherent processing stages. Medium complexity selects 6–8 stages.
Module roles: Infrastructure (models, base class, config, exceptions, registry), stages (processing steps implementing a common ABC), adapters (wrapping stages with additional behavior), middleware (cross-cutting decorators), utilities, and legacy/distractor modules.
Anti-triviality: (1) Registry wiring—stages connected via config, not direct imports; (2) Adapter indirection—ABC interface layer; (3) Distractor modules—files not in the main pipeline; (4) Neutral naming—mod_a.py, not extract.py; (5) Hidden invariants—constraints discoverable only by reading function bodies or tests.
4.2 Four Edge Types
Each generated codebase contains edges of four types, reflecting different methods of discovery:
imports (67%): Python import statements. Discoverable by AST parsing.
calls_api (17%): Runtime function calls between modules. Requires reading function bodies.
registry_wires (9%): Config-driven connections where a registry module dynamically loads stages via importlib based on names listed in a JSON config file. No static import statement exists. Requires reading both the config and the registry’s loading logic.
data_flows_to (7%): Data dependency where one module’s output feeds another’s input. Requires understanding orchestration logic.
These proportions are not tunable parameters but emerge from the architectural template: a Pipeline with stages naturally generates 3–4 more imports edges because every module requires static imports, while semantic edge types arise only at specific architectural boundaries (adapterstage, orchestratorstage, consecutive stages). This mirrors real-world Pipeline codebases, where import edges dominate. Different architectural patterns (event-driven, microservices) would yield different distributions—a key motivation for extending the generator. Crucially, roughly one-third of edges are invisible to import-following (specific to our Pipeline template; other architectural patterns would yield different proportions), creating a meaningful gap between syntactic analysis and semantic understanding.
4.3 Invariant Planting
Each codebase contains 15–16 planted constraints across five types: boundary (forbidden dependencies), dataflow (required processing chains), interface (access-only-through-ABC), invariant (naming/structural conventions), and purpose (design rationales). Every constraint has a structured canonical form (type, src, dst, via, pattern) for machine-comparable scoring and at least one evidence source (test file, structural pattern, or documentation).
4.4 Generation Statistics
| Metric | Seed 42 | Seed 123 | Seed 999 |
| Modules | 27 | 30 | 27 |
| Total edges | 70 | 84 | 70 |
| imports | 47 (67%) | 56 (67%) | 47 (67%) |
| calls_api | 12 (17%) | 15 (18%) | 12 (17%) |
| data_flows_to | 5 (7%) | 6 (7%) | 5 (7%) |
| registry_wires | 6 (9%) | 7 (8%) | 6 (9%) |
| Invariants | 15 | 16 | 15 |
| Sub-packages | 5 | 5 | 5 |
5 Metrics
Dependency F1.
Compare predicted edges from the cognitive map against ground truth. An edge matches if (source, target, type) all agree. We report precision, recall, and F1 separately.
Invariant F1.
Match agent-discovered constraints against planted constraints via structured form comparison—exact field matching on (type, src, dst, via), not text similarity. This eliminates ambiguity in natural-language descriptions.
Confidence Calibration.
Expected Calibration Error (ECE) between agent-stated per-edge confidence and actual correctness, binned into 5 confidence intervals.
Two Efficiency Curves.
Action-efficiency: AUC(F1 vs. total actions)—the headline metric. Observation-efficiency: AUC(F1 vs. OPEN count)—diagnostic, isolating information gain per file opened. Belief is piecewise-constant between probes; integration is trapezoidal.
Active-Passive Gap.
For each metric : , decomposed into selection and decision components as described in §3.
Belief Revision Score (BRS).
After a mutation at time : . Decomposed into Inertia-proper (correctly-believed elements not updated after evidence) and Impact-discovery (missing elements newly found). A sham condition (evidence without actual change) measures Gullibility. (BRS evaluation is implemented but reserved for future work.)
6 Experiments
We evaluate four rule-based exploration strategies and five frontier LLM agents from three providers on three generated codebases (seeds 42, 123, 999) under identical conditions: budget , probe interval . These results are preliminary; we report them to validate the benchmark’s discriminative power and invite community replication.
6.1 Methods
Rule-based baselines.
Oracle: Outputs the ground-truth graph directly. Upper bound (F1 = 1.0).
Config-Aware: Lists all directories, opens config/registry files first, parses module references, then follows imports BFS.
Random: Lists root, then opens files uniformly at random until budget exhausted. Builds map from observed imports.
BFS-Import: Opens files breadth-first following import chains.
All baselines build cognitive maps from AST-parsed imports and config references.
LLM agents.
Claude Sonnet 4.6: Anthropic’s frontier coding model (claude-sonnet-4-6).
GPT-5.3-Codex: OpenAI’s code-specialized reasoning model (gpt-5.3-codex). A reasoning model (temperature fixed at 1 per API constraint).
Gemini 2.5 Flash: Google’s reasoning-optimized flash model (gemini-2.5-flash). A “thinking” model with extended internal reasoning.
Gemini 3.1 Pro: Google’s frontier pro model (gemini-3.1-pro-preview).
Gemini 3 Flash: Google’s frontier flash model (gemini-3-flash-preview).
All models receive identical prompts: a system prompt describing the partial observability environment and available actions, per-turn action prompts showing remaining budget and opened files, and structured JSON probe prompts with schema and example (Appendix A). Temperature is set to 0 except for GPT-5.3-Codex (reasoning model constraint). All probe responses were parsed successfully via the deterministic repair parser (zero format failures across all runs).
6.2 Results
| Method | Type | Dep F1 | Precision | Recall | AUC | Inv F1 | Files |
| Oracle | baseline | 1.000 | 1.000 | 1.000 | — | 0.000 | — |
| Config-Aware | baseline | 0.577 | 0.736 | 0.475 | 0.212 | 0.000 | 13 |
| Random | baseline | 0.538 | 1.000 | 0.368 | 0.142 | 0.000 | 13 |
| BFS-Import | baseline | 0.293 | 1.000 | 0.173 | 0.079 | 0.000 | 13 |
| Claude Sonnet 4.6 | LLM | 0.646 | 0.991 | 0.479 | 0.377 | 0.000 | 12 |
| GPT-5.3-Codex | LLM | 0.564 | 0.768 | 0.460 | 0.243 | 0.000 | 14 |
| Gemini 2.5 Flash | LLM | 0.328 | 0.568 | 0.232 | 0.227 | 0.000 | 13 |
| Gemini 3.1 Pro | LLM | 0.262 | 0.894 | 0.156 | 0.135 | 0.000 | 9 |
| Gemini 3 Flash | LLM | 0.129 | 0.713 | 0.071 | 0.120 | 0.000 | 12 |
| Method | imports | calls_api | data_flows | reg._wires |
| Config-Aware | 0.58 | 0.00 | 0.00 | 1.00 |
| Random | 0.55 | 0.00 | 0.00 | 0.00 |
| BFS-Import | 0.25 | 0.00 | 0.00 | 0.00 |
| Claude Sonnet 4.6 | 0.55 | 0.15 | 0.00 | 1.00 |
| GPT-5.3-Codex | 0.57 | 0.15 | 0.31 | 0.32 |
| Gemini 2.5 Flash | 0.32 | 0.10 | 0.00 | 0.00 |
| Gemini 3.1 Pro | 0.20 | 0.05 | 0.06 | 0.05 |
| Gemini 3 Flash | 0.09 | 0.08 | 0.00 | 0.00 |
6.3 Analysis
Claude Sonnet 4.6 surpasses baselines.
Claude Sonnet 4.6 achieves F1 = 0.646, surpassing the best rule-based strategy (Config-Aware, 0.577) by 7 points—the first LLM agent to outperform all baselines on this benchmark. It opens 12 files (vs. 13 for baselines) yet achieves higher recall (0.479 vs. 0.475), with near-perfect precision (0.991: only 1 false positive across all 3 codebases). Critically, it discovers all 19 ground-truth REGISTRY_WIRES edges with zero false positives (Table 3)—matching Config-Aware’s recall but without its 38 false positives (Config-Aware achieves only 33% precision on this edge type). This suggests Claude reads both the config file and registry loading logic, correctly inferring which modules are dynamically wired.
LLM agents discover all four edge types.
Table 3 reveals the most distinctive finding: LLM agents collectively discover all four edge types, while baselines find at most two (IMPORTS and REGISTRY_WIRES for Config-Aware). All five LLMs found CALLS_API edges—invisible to all rule-based baselines—confirming genuine code comprehension beyond syntactic parsing. GPT-5.3-Codex is the only method to discover DATA_FLOWS_TO edges (5 of 16, 31% recall, with perfect precision), demonstrating understanding of how data flows through the pipeline orchestrator. These edges require multi-hop reasoning: tracing one module’s return value through the runner into the next module’s input.
GPT-5.3-Codex: strong but variable.
GPT-5.3-Codex achieves F1 = 0.564 on average (second-highest), but with high variance across codebases: 0.719 on seed 42 vs. 0.392 on seed 999. It discovers the broadest range of edge types (all four), including 5 DATA_FLOWS_TO edges with perfect precision. However, it over-generates REGISTRY_WIRES edges (6 correct, 29 false positives across 3 codebases), reducing its overall precision to 0.768. The high variance may stem from sensitivity to early exploration choices: different initial file sequences appear to trigger different architectural hypotheses, with some runs “locking in” to a more productive exploration path.
The belief externalization bottleneck.
The three Gemini models illustrate the belief externalization problem most clearly. Gemini 2.5 Flash opens 13 files (comparable to baselines) but recalled only 32% of IMPORTS edges, versus 55–58% for Config-Aware and Random. It appears to understand more than it reports: the model follows import chains when choosing exploration actions (evidence from action logs) but fails to transcribe observed imports into the cognitive map JSON. Gemini 3 Flash exhibits an extreme version: opening 12 files but reporting only 7 edges on average, with one codebase yielding just 4 edges total. This suggests that belief externalization—faithfully serializing internal understanding into a structured JSON schema—is itself a non-trivial capability. The gap between what an agent knows and what it writes down is a first-order confounder in any belief-probing benchmark.
Gemini 3.1 Pro: conservative but under-explored.
Gemini 3.1 Pro opens only 9 files on average (vs. 12–14 for other models), appearing to terminate exploration prematurely or allocate budget to LIST/SEARCH actions rather than OPEN. Despite high precision (0.894), its low file coverage directly limits recall (0.156). This suggests that the model’s exploration strategy—not its comprehension ability—is the primary bottleneck.
Precision-recall profiles differ qualitatively.
Baselines using AST parsing (BFS-Import, Random) achieve perfect precision—every reported edge is correct. Among LLMs, Claude Sonnet 4.6 also achieves near-perfect precision (0.991) while attaining the highest recall, suggesting that strong models can externalize beliefs both accurately and comprehensively. In contrast, Gemini 2.5 Flash has the lowest precision (0.568), generating 19 false CALLS_API edges and 8 false DATA_FLOWS_TO edges alongside its correct discoveries. Config-Aware shows a parallel pattern: perfect IMPORTS precision but only 33% REGISTRY_WIRES precision due to aggressive config-based inference.
Efficiency strongly favors LLM agents.
Under the Action AUC metric—which rewards early discovery—Claude Sonnet 4.6 dominates (0.377), nearly doubling Config-Aware (0.212). All five LLMs except Gemini 3 Flash (0.120) outperform BFS-Import (0.079) on AUC. This reflects the models’ ability to form useful architectural hypotheses from the first few files, while rule-based methods require more observations before deterministic parsing yields comparable coverage.
Invariant discovery is uniformly zero.
All methods score Invariant F1 = 0.0. LLM agents do report invariants (Claude averages 8.0, Gemini 2.5 Flash 6.0, GPT-5.3-Codex 5.7 per codebase), but none match the planted ground-truth constraints under structured form comparison. The Oracle baseline also scores zero: it provides dependency edges only, and invariant discovery requires natural-language comprehension of test assertions and documentation that no current method attempts.
Random’s competitiveness and budget effects.
Random exploration (F1 = 0.538) performs nearly as well as Config-Aware (0.577). This near-parity reflects the current scale: budget on 27–30 modules allows opening roughly half the files, making raw coverage a strong proxy for import discovery. A budget sweep (Appendix C) confirms that at , Config-Aware’s advantage grows to 3 over Random (0.175 vs. 0.056). The meaningful discrimination at lies in understanding what you’ve seen: LLM agents discover qualitatively different edge types despite comparable file coverage.
7 Discussion
7.1 Implications for Code Agent Design
Current code agents (SWE-agent, Aider, Claude Code, Cursor) use tool-augmented retrieval to navigate repositories, but none explicitly construct or maintain a structured architectural belief. Our results show that the strongest LLM agents can surpass rule-based strategies when they combine semantic comprehension with reliable belief externalization (Claude Sonnet 4.6), but that most models fail at the externalization step despite demonstrating code understanding in their exploration behavior.
This suggests three improvement paths: (1) hybrid approaches combining AST-level import extraction with LLM semantic analysis, ensuring structural completeness while adding semantic depth; (2) belief externalization training—explicitly optimizing models to faithfully serialize architectural knowledge into structured formats; and (3) exploration strategy optimization—models like Gemini 3.1 Pro that under-explore (opening only 9 files) could improve significantly with better budget allocation.
7.2 The Belief Externalization Problem
The spread across models is striking: Claude Sonnet 4.6 recalled 55% of IMPORTS edges with 100% precision, while Gemini 3 Flash recalled only 9% despite opening a similar number of files. This 6 gap in IMPORTS recall—for edges the models demonstrably saw—suggests that belief externalization capability varies dramatically across model families, not just between LLMs and baselines.
Interestingly, Claude Sonnet 4.6’s near-perfect precision (0.991) across all edge types shows that faithful externalization is achievable with current models. The challenge appears model-specific rather than fundamental: it is not that structured JSON is inherently difficult to produce, but that some models have not been optimized for this capability. Future work should investigate whether fine-tuning, chain-of-thought prompting, or iterative self-checking can close this gap for weaker models.
7.3 Limitations
ToCS v0.1 is deliberately narrow in scope. Generated codebases, while structurally realistic, may not capture all patterns of real-world architecture. In particular, our use of neutral filenames (mod_a.py rather than authentication_service.py) removes naming priors that real codebases provide, making the task harder in one dimension while keeping it easier in others (clean structure, no organic complexity, limited dead code). This design choice isolates architectural reasoning from lexical guessing, but means our difficulty estimates may not transfer directly to real repositories. We evaluate a single architectural pattern (Pipeline) in a single language (Python). Our LLM results are preliminary: five models from three providers, each evaluated on 3 codebases with a single run per model-codebase pair (no variance estimation across runs). GPT-5.3-Codex’s high cross-codebase variance (F1 ranging from 0.392 to 0.719) highlights the need for repeated runs. The cognitive map externalization is a proxy for internal representation—a model may “know” more than it serializes, as the wide variation across models suggests. We plan to validate via correlation between map quality and counterfactual probe accuracy, and to expand to additional models (open-weight), architectural patterns, repeated runs with variance estimation, and the REVISE evaluation mode.
7.4 Community Call
We release the complete ToCS toolkit—codebase generator, partial observability harness, evaluation pipeline, and scoring metrics—as open-source software. We invite the community to:
1.
Evaluate additional models: Run open-weight models (Llama, Qwen, DeepSeek) and replicate our frontier model results with repeated runs.
2.
Contribute codebase patterns: Extend beyond Pipeline to event-driven, microservices, plugin systems, hexagonal architecture (see CONTRIBUTING.md).
3.
Add languages: The harness is language-agnostic; the generator needs new templates for TypeScript, Go, etc.
4.
Test scaffold-augmented agents: Evaluate whether static analysis tools, RepoMap-style context, or explicit state tracking improve map quality.
8 Conclusion
We introduced ToCS, a benchmark for evaluating whether code agents can construct coherent architectural beliefs through active codebase exploration under partial observability. Our procedural generator produces medium-complexity codebases with four typed edge categories and planted invariants. Preliminary experiments with four baselines and five frontier LLM agents from three providers reveal that: (1) exploration strategy limits discovery for rule-based methods, but the strongest LLM agent (Claude Sonnet 4.6, F1 = 0.646) surpasses all baselines by combining semantic comprehension with reliable belief externalization; (2) LLM agents collectively discover all four edge types—including DATA_FLOWS_TO edges (GPT-5.3-Codex, 31% recall) and REGISTRY_WIRES edges (Claude Sonnet 4.6, 100% recall)—capabilities entirely absent from rule-based methods; yet (3) weaker LLM agents score below simple heuristics due to severe belief under-reporting, with a 6 spread in IMPORTS recall across models (9–57%), confirming that structured belief externalization remains a key capability gap. The 35-point F1 gap between Oracle and the best LLM agent, combined with zero invariant discovery across all methods, defines a challenging and discriminative evaluation space. We release ToCS as an open benchmark and invite the community to replicate and extend these preliminary results.
Appendix A Evaluation Prompts
System prompt (excerpt).
The agent is told: “You are an expert software engineer exploring a codebase you have never seen before. You are operating under PARTIAL OBSERVABILITY […] Available actions: LIST, OPEN, SEARCH, INSPECT, DONE. […] Your goal is to build a complete and accurate ARCHITECTURAL UNDERSTANDING.” The full prompt describes action semantics and exploration strategy tips.
Cognitive map probe.
After every actions, the agent receives: “Externalize your current architectural belief as JSON matching the following schema […]” The schema specifies components with status/purpose/exports/edges/confidence, invariants with structured canonical form, and an uncertainty summary. A worked example demonstrates the expected format (see supplementary materials).
Appendix B Generated Codebase Example
A medium-complexity codebase (seed 42, text processing domain) generates the following structure:
text_processor/ +-- __init__.py +-- middleware/ +-- models.py | +-- mod_i.py (logging) +-- base.py | +-- mod_j.py (retry) +-- config.py +-- utils/ +-- exceptions.py | +-- helpers.py +-- registry.py | +-- formatters.py +-- stages/ +-- legacy/ | +-- mod_a.py | +-- old_pipeline.py | +-- mod_b.py | +-- compat.py | +-- mod_c.py +-- runner.py | +-- mod_d.py +-- cli.py | +-- mod_e.py +-- pipeline_config.json | +-- mod_f.py +-- test_smoke.py +-- adapters/ | +-- mod_g.py | +-- mod_h.py
Stages implement a StageBase ABC. The registry loads stage ordering from pipeline_config.json. Stages cannot import each other directly (enforced by test). Legacy modules are distractors not in the pipeline flow.
Appendix C Budget Sweep
To characterize how exploration budget affects discrimination between strategies, we ran all four baselines at across 3 codebases. Table 4 reports mean Dependency F1.
| Method | ||||
| Oracle | 1.000 | 1.000 | 1.000 | 1.000 |
| Config-Aware | 0.175 | 0.492 | 0.577 | 0.626 |
| Random | 0.056 | 0.317 | 0.538 | 0.632 |
| BFS-Import | 0.078 | 0.157 | 0.293 | 0.603 |
At tight budgets (), Config-Aware’s strategy of prioritizing registry files yields a 3 advantage over Random (0.175 vs. 0.056). As budget increases, raw coverage dominates: at (allowing agents to open nearly all files), Random matches Config-Aware (0.632 vs. 0.626). BFS-Import shows the steepest growth from to (0.293 0.603), reflecting its systematic but slow import-chain traversal. The convergence at high budgets confirms that the current medium-complexity codebases require tighter budgets or larger codebases to strongly discriminate exploration strategy for structural edge types; the meaningful discrimination at lies in semantic edge discovery by LLM agents.